OpenStudy (anonymous):

plez help For the linear equation, x + y - 1 = 0, determine a corresponding value of y for the x-value of -6. Type a numerical answer

OpenStudy (triciaal):

x + y = 1 y = 1 - x when x = -6 then y = 1 - (-6) = 7